40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 5 09:20:44 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_Chamber_optical_layout.pdf
    Reply  Thu Aug 5 12:39:56 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_in-air_setup.pdf
       Reply  Thu Aug 5 14:29:42 2021, Stephen, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber IMG_9685.JPGIMG_9687.JPGrouting_markup_of_IMG_9688.pngIMG_9682.JPG
          Reply  Tue Aug 24 15:01:25 2021, Stephen, Laser, 2um Photodiodes, 2um PD in new cryo chamber 
             Reply  Thu Aug 26 14:31:46 2021, Aidan, Stephen, Laser, 2um Photodiodes, Incident power calibration of 2um PD in new cryo chamber IMG_4420.jpgREF_PD_POW_CAL_cryo.pdf
                Reply  Fri Aug 27 17:57:30 2021, Stephen, DailyProgress, 2um Photodiodes, Start of PD pumpdown  
                   Reply  Sat Aug 28 07:01:21 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  20210828_-_IMG_4445.jpg20210828_-_IMG_4443.jpg
                      Reply  Sun Aug 29 11:17:43 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  Screen_Shot_2021-08-29_at_11.16.55_AM.pngScreen_Shot_2021-08-29_at_11.22.03_AM.png
                         Reply  Mon Aug 30 08:04:52 2021, Aidan, DailyProgress, 2um Photodiodes, Monday morning observation of JPL A1 PD cool down  Screen_Shot_2021-08-30_at_8.04.15_AM.png
Message ID: 2655     Entry time: Sat Aug 28 07:01:21 2021     In reply to: 2654     Reply to this: 2656
Author: Aidan 
Type: DailyProgress 
Category: 2um Photodiodes 
Subject: Start of PD pumpdown  

Note on the script - it's running in a TMUX session on WS1: On for 15s, Off for 15s. A much better way to do this would've been to turn on an EXC in the laser current channel (I only remembered the AWG once I'd left). I'll pop in this morning to check on things and see if I can change the modulation.

The purpuose of the laser modulation is to continually monitor the PD dark current level as well as the change in the nominal QE as a function of temperature. I might change it to a sawtooth wave - if the PD is saturating from too much intensity, we'll see a decrease in the QE each time the laser power increases. FYI - we're getting about 0.8mW onto the PD and it should be close to the beam waist (numbers to follow on that); 

The DAC outputs are really wandering around a lot. We should diagnose what's going on there as this makes the laser current set point very noisy. 

10:00AM Saturday update: Temperature of the workpiece (photodiode) is around 192K. I terminated the script and started an excitation RAMP on the laser diode current. The response of the diode has increased dramatically and I think I see signs of the QE rolling off as the incident power increases.

Quote:

[Aidan, Stephen]

After Aidan validated that model inputs were creating physical parameter changes, we proceeded with some last few checks before closing up. Notes:

 - Aidan set up a helpful script turning laser power on and off, and strip tool to follow PD and monitor PD signals. A strip tool chart was used to confirm that there was no loss of functionality or alignment during pump down.

 - Checked Heater function and Workpiece RTD response - all good.

 - Confirmed alignment by steering at periscope output mirror and watching PD voltage.

At this point Aidan gave a green light for pumpdown prep, and to start pumpdown and start cooldown. Notes:

 - Stephen disconnected the PD connector accidentally while trying to add strain relief (the irony is palpable). Reattachment of connector didn't seem to affect any signal levels.

 - During installation of radiation shield lids, shields became misaligned and PD signal fell (presumably due to clipping). Recovered previous signal levels by realligning outer shield.

 - Double checked that everything seemed good to go, no issues!

 - Timeline of pumpdown and cooldown:

     17:32 - Pumpdown started.

     17:47 - Turbo spin up started (pump station delay parameter).

     17:55 - Pressure dropped below 1 mTorr, so I started cryocooler.

     18:05 - Healthy so far - Pressure had come down half a decade more, and Coldhead RTD was reading 235 K.

 

Attachment 1: 20210828_-_IMG_4445.jpg  5.412 MB  Uploaded Sat Aug 28 12:39:04 2021  | Hide | Hide all
20210828_-_IMG_4445.jpg
Attachment 2: 20210828_-_IMG_4443.jpg  2.603 MB  Uploaded Sat Aug 28 12:39:14 2021  | Hide | Hide all
20210828_-_IMG_4443.jpg
ELOG V3.1.3-